OpenStudy (anonymous):

Help ASAP Trevor rode 3/16 of a mile every minute. How many minutes did it take him to ride 20 1/2 miles? Round your answer to the nearest minute.

OpenStudy (kropot72):

You need to find how many 3-sixteenths are in 20 1/2 minutes. \[\large time=\frac{\frac{41}{2}}{\frac{3}{16}}=\frac{41\times16}{2\times3}=you\ can\ calculate\]

OpenStudy (anonymous):

he road 3/16 of a mile every minute. you need to know how many minutes it took him to ride 20 1/2? first change 20 1/2 into an improper fraction over 16 like 3/16 is. so it is now 328/16 now you need to divide: 328/16 divided by 3/16.... which will give you the minutes